# Advances on the finite element discretization of fluid-structure interaction problems

### v1

@@ Abstract.
We review the main features of an unfitted finite element method for interface and fluid-structure interaction problems based on a distributed Lagrange multiplier in the spirit of the fictitious domain approach. We recall our theoretical findings concerning well-posedeness, stability, and convergence of the numerical schemes, and discuss the related computational challenges. In the case of elliptic interface problems, we also present a posteriori error estimates.
@@ Introduction
The mathematical description of many physical and engineering problems requires partial differential equations involving different operators or discontinuous coefficients so that the computational domain is partitioned into several (possibly time dependent) regions. The coupling between different models is enforced at the interface via suitable transmission conditions. Among the applications, we mention, for instance, thermo-mechanical problems, cardiac simulations, computational geo-science, and many others.
@@
When designing effective numerical methods for the mentioned problems, an accurate representation of the underlying geometry is needed. Fitted approaches are based on a mesh which conforms to the interface and provide accurate solutions [11, 56, 57, 63, 35]. However, when applied to evolution problems, computational issues may arise. For instance, the Arbitrary Lagrangian–Eulerian approach [54, 42, 59], appealing for fluid-structure interaction problems, needs an update of the mesh at each time step. In case of large displacements or deformations, the update may break shape regularity, thus requiring the generation of a new mesh, a costly operation in general. Moreover, when fluid and solid have similar or equal densities, the added mass effect may affect the performance of the method [33]. Such phenomenon may lead to the unconditional instability of the scheme regardless the physical parameters. The added mass effect can be reduced by an appropriate treatment of transmission conditions, as described in [41, 9].
@@
Due to the critical aspects of fitted approaches, several unfitted methods have been introduced during the past decades. In this case, meshes are generated without taking care of the interface, which is allowed to cross elements. We mention, for instance, the Immersed Boundary Method [66, 67], where the interface position is tracked by Dirac delta functions, and the level set method [71, 34], where the interface corresponds to the zero level set of a certain function. Other techniques such as Nitsche-XFEM [1], Cut-FEM [28, 31, 52], Finite Cell methods [38, 39, 48], and Ghost-FEM [7] are based on the use of penalty terms. The fictitious domain approach has been introduced to model the particulate flow as described in [50, 49], while the application to more general cases can be found in [75, 14].
@@
In [8, 17], we introduced a new fictitious domain formulation with distributed Lagrange multiplier (FD-DLM) for the discretization of elliptic interface and fluid-structure interaction problems. Indeed, these problems share the common feature that the domain is divided into two regions separated by an interface. In the case of elliptic interface problems this is due to the presence of discontinuous coefficients, while fluid-structure interactions involve different types of equations, corresponding to constitutive models for fluid and solid. This survey paper aims to collect in a single place the main properties of this approach, including the well-posedness and stability of the solution, convergence of the discrete scheme and computational aspects.
@@
The main idea of our approach is to extend one domain into the other. The two domains are then discretized by fixed independent meshes. In particular, in case of fluid-structure interactions, the fluid dynamics is described in an Eulerian framework on the extended domain, while the deformation of the solid body is represented in a Lagrangian setting by considering a reference domain which is mapped, at each time instant, into the actual configuration. In order to impose that the solution corresponding to the fictitious region coincides with the one of the immersed domain, we consider a Lagrange multiplier and a coupling term that are added to the model. We emphasize that all computations are done on fixed domains on meshes that are generated only once.
@@
The paper is organized into two parts. In Section 1, we focus on elliptic interface problems and discuss in some details the main features of our approach, which will be extended later to fluid-structure interactions. After presenting the continuous and discrete problems, we describe how to deal with the coupling term [18, 19]. The implementation of its finite element approximation requires integration over non-matching grids, which is a challenging aspect of several immersed methods [60, 46, 62, 27, 47]. Next, we discuss the design of effective block preconditioners, whose action is performed through direct inversion [22] or by employing multigrid algorithms [63]. Another issue is related to the fact that the solution of the interface problem can present singularities along the interface. A way to achieve optimal convergence properties is to employ adaptive mesh refinement based on suitable error indicators. This crucial aspect has been previously analyzed in e.g. [13, 37, 69, 32] for other numerical schemes. Here we report the a posteriori error analysis presented in [4] for our FD-DLM formulation.
@@
In Section 2, we present the extension of the FD-DLM formulation and the related techniques to fluid-structure interaction problems, where an incompressible structure made of viscous hyper-elastic material is immersed in an incompressible Newtonian fluid (see [17] and references therein). In this case, the regions occupied by the fluid and the solid evolve in time and their configuration is itself an unknown of the problem. The actual position of the solid is obtained by mapping the reference domain through the deformation unknown: this gives further difficulties for the computation of the so-called coupling terms. Viscous hyper-elastic materials are useful when dealing with several applications: we mention, for instance, blood flow in heart [66, 68] and vessels [40, 55, 10], and flapping flexible filaments in soap films [76, 77]. This choice of solid constitutive law is relevant for our formulation since it allows to separate the viscous and elastic parts of the Cauchy stress tensor so that we can treat each of them in the appropriate Eulerian and Lagrangian frameworks, respectively.
@@
Both sections 1 and 2 are completed by some numerical results that confirm the theoretical findings and show the versatility of the proposed method.
@@
Sobolev spaces are denoted by the standard symbol $W^{s,p}(D)$, where $s\in\mathbb{R}$ refers to differentiability and $p\in[1,\infty]$ is the integrability exponent. If $p=2$, then the notation $H^{s}(D)=W^{s,2}(D)$ is employed, with associated norm $\|\cdot\|_{s,D}$ and semi-norm $|\cdot|_{s,D}$. The $L^{2}i(D)$ scalar product is denoted as usual by $(\cdot,\cdot)_{D}$.
@@ 1. Fictitious domain formulation for elliptic interface problems
Let $\Omega\subset\mathbb{R}^{d}$, with $d=2,3$, be an open bounded domain with Lipschitz boundary $\partial\Omega$. The domain $\Omega$ is made of two disjoint regions, $\Omega_{1}$ and $\Omega_{2}$, separated by the interface ${\Gamma=\overline{\Omega}_{1}\cap\overline{\Omega}_{2}}$. We assume $\Gamma$ to be Lipschitz continuous. Two examples of domain configuration for $d=2$ are sketched in Figure 1.
@@
$-\operatorname{div}\left(\nu_{1}\operatorname{\nabla}u_{1}\right)$ $=f_{1}$ in $\text{in }\Omega_{1}$
$-\operatorname{div}\left(\nu_{2}\operatorname{\nabla}u_{2}\right)$ $=f_{2}$ in $\text{in }\Omega_{2}$
$u_{1}-u_{2}$ $=0$ on $\text{on }\Gamma$
$\nu_{1}\,\nabla u_{1}\cdot\mathbf{n}_{1}+\nu_{2}\,\nabla u_{2}\cdot\mathbf{n}_{2}$ $=0$ on $\text{on }\Gamma$
$u_{1}$ $=0$ on $\text{on }\partial\Omega_{1}\setminus\Gamma$
$u_{2}$ $=0$ on $\text{on }\partial\Omega_{2}\setminus\Gamma.$
@@
$\tilde{\nu}=\begin{cases}\nu_{1}&\text{ in }\Omega_{1}\\ \nu_{2}&\text{ in }\Omega_{2},\end{cases}\qquad\tilde{f}=\begin{cases}f_{1}&\text{ in }\Omega_{1}\\ f_{2}&\text{ in }\Omega_{2},\end{cases}$
the variational formulation of Problem 1 can be written by applying standard arguments as follows:
@@
Equation (1) is equivalent to Problem 1 with $\tilde{u}_{|\Omega_{i}}=u_{i}$ with $i=1,2$.
Remark 1.
In general, due to the jump of the normal derivative of $\tilde{u}$ on $\Gamma$, the solution of (1) does not belong to $H^{2}(\Omega)$. Indeed, it is well-known that the regularity of the solution to elliptic interface problems with discontinuous coefficients on a domain with Lipschitz boundary belongs to $H^{r}(\Omega)$ with $21<r<3/$ 2. However, it is possible to show that the restriction $\tilde{u}_{|\Omega_{i}}$, $i=1,2$, is more regular depending on the smoothness of $\Gamma$. If $\Gamma$ is only Lipschitz continuous and presents re-entrant corners, then $u_{i}\in H^{s}(\Omega_{i})$ with $3/2<s\leq 2$ (see [64]).
From Remark 1, solving (1) by the finite element method on a mesh which does not take into account the presence of $\Gamma$, gives a non-optimal order of convergence. An optimal method can be recovered by constructing a mesh fitted with the interface $\Gamma$. Here by optimal we mean full rate of convergence for the polynomial degree of the finite element space. Clearly, this requirement is strong, especially if the problem under consideration originates from the time discretization of a more complex one, such as a fluid-structure interaction problem. In such case, the interface $\Gamma$ may evolve in time and, thus the mesh should be updated according to the position of $\Gamma$ at each time step.
The method we proposed is based on a fictitious domain formulation and a distributed Lagrange multiplier. Our method is not optimal in the sense explained above, but it allows for a natural extension to moving interfaces.
@@ 1.1. Fictitious domain formulation
We denote by $u\in H^{1}_{0}(\Omega)$ the extension of $u_{1}$ to $\Omega$ satisfying that $u_{|\Omega_{1}}=u_{1}$. The extended $u$ is required to match $u_{2}\in H^{1}(\Omega_{2})$ in $\Omega_{2}$. We enforce this superimposition between $u_{|\Omega_{2}}$ and $u_{2}$ at variational level by introducing a suitable functional space $\Lambda$ and a bilinear form $c:\Lambda\times H^{1}(\Omega_{2})\rightarrow\mathbb{R}$ satisfying
@@
[…]
with $\|\varphi\|_{1,\Omega_{2}}=\|\mu\|_{\Lambda}$, but give rise to different discretizations.
We introduce the Lagrange multiplier $\lambda\in\Lambda$ associated with the condition $u_{|\Omega_{2}}=u_{2}$, then the weak formulation of Problem 1 in a fictitious domain setting is:
Problem 2.
@@
In [8, Thm. 2], it was proved that Problem 2 is equivalent to the standard variational formulation in (1). The following proposition states the well-posedness of Problem 2, see [8, Prop. 1].
Proposition 1.
Let $\nu$ and $\nu_{2}$ be two positive constants. Given $f\in L^{2}(\Omega)$ and $f_{2}\in L^{2}(\Omega_{2})$, Problem 2 admits a unique solution ${(u,u_{2},\lambda)\in H^{1}_{0}(\Omega)\times H^{1}(\Omega_{2})\times\Lambda}$ satisfying
@@
Since the problem has a saddle point structure, its well-posedness is established by proving the following sufficient conditions [15]:
@@
Remark 2.
The stability estimate in Proposition 1 depends on the choice of $\Lambda$ and $c(\cdot,\cdot)$ through the constant $C$.
Remark 3.
Notice that, due to the equivalence of Problem 2 and (1), we have that the extended solution $u$ belongs to $H^{r}(\Omega)$ with $21<r<3/$ 2. On the other hand, $u_{2}$ belongs to $H^{s}(\Omega_{2})$ with $3/2<s\leq 2$, see Remark 1.
@@ 1.2. Discretization
We discretize Problem 2 by mixed finite elements. We partition $\Omega$ and $\Omega_{2}$ by two independent shape regular meshes $\mathcal{T}_{h}$ and $\mathcal{T}_{h}^{2}$, respectively. We denote by $h$ the size of $\mathcal{T}_{h}$, whereas $h_{2}$ denotes the mesh size of $\mathcal{T}_{h}^{2}$. The arguments we are going to present in this section are valid for both triangular and quadrilateral meshes, although we present in detail the quadrilateral case.
@@
On the other hand, if $\Lambda=H^{1}(\Omega_{2})$, the inner product in $H^{1}(\Omega_{2})$ is used also at discrete level.
The discrete formulation of Problem 2 reads as follows.
Problem 3.
@@
For a generic element $K$, we denote by $\mathcal{P}_{1}(K)$ the space of linear polynomials if $K$ is a simplex and by $\mathcal{Q}_{1}(K)$ the space of bilinear polynomials if $K$ is a quadrilateral or an hexahedron. Moreover, let $\mathcal{P}_{0}(K)$ be the space of constants and $\mathfrak{B}(K)$ the space of biquadratic bubbles vanishing at the boundary of an element $K$.
We solve Problem 3 by considering two distinct sets of finite element spaces, which are distinguished by the Lagrange multiplier space being continuous or discontinuous depending on the choice of the coupling term. More precisely, we consider continuous linear elements for all variables [8], that is
(6) $V_{h}=\left\{v_{h}\in V:v_{h|K}\in\mathcal{Q}_{1}(K),\forall K\in\mathcal{T}_{h}\right\},$
$V_{2,h}=\left\{v_{2,h}\in V_{2}:v_{2,h|K_{2}}\in\mathcal{Q}_{1}(K_{2}),\forall K_{2}\in\mathcal{T}_{h}^{2}\right\},$
$\Lambda_{h}=V_{2,h}$
@@
(7) $V_{h}=\left\{v_{h}\in V:v_{h|K}\in\mathcal{Q}_{1}(K),\forall K\in\mathcal{T}_{h}\right\}\subset V,$
$V_{2,h}=\left\{v_{2,h}\in V_{2}:v_{2,h|K_{2}}\in\mathcal{Q}_{1}(K_{2})\oplus\mathfrak{B}(K_{2}),\forall K_{2}\in\mathcal{T}_{h}^{2}\right\}\subset V_{2},$
$\Lambda_{h}=\left\{\lambda_{h}\in\Lambda:\lambda_{h|K_{2}}\in\mathcal{P}_{0}(K_{2}),\forall K_{2}\in\mathcal{T}_{h}^{2}\right\}\subset\Lambda.$
@@
Both choices are stable since they satisfy the discrete counterpart of (3) and (4). In particular, stability of (6) was proved for both simplicial and quadrilateral/hexahedral meshes in [8, 25]. The ellipticity in the kernel condition is met under the assumption that $\nu_{2}>\nu$. This requirement can be relaxed for simplicial meshes if $h_{2}/h^{d}$ is small enough, see [25]. Stability of (7) was proved in [3]. The ellipticity in the kernel condition is satisfied assuming that $\nu_{2}>\nu$. Numerical findings suggest that this requirement might be relaxed [3].
The error estimates of the discrete solution can be found in [8, 25, 3]. We point out that such bounds depend on the choice of $\Lambda_{h}$. The following estimate holds.
Proposition 2.
Let $(u,u_{2},\lambda)$ and $(u_{h},u_{2h},\lambda_{h})$ be the solutions of Problems 2 and 3, respectively. If the discrete spaces are given by (6), then the following error estimate holds true
@@
where $r$, $s$ are given in Remark 1, $11/2<t<$ 1 and $C_{1},\,C_{2}$ are positive constants independent of $h$ and $h_{2}$.

### v2

# Advances on finite element discretization of fluid-structure interaction problems

@@ Abstract.
We review the main features of an unfitted finite element method for interface and fluid-structure interaction problems based on a distributed Lagrange multiplier in the spirit of the fictitious domain approach. We recall our theoretical findings concerning well-posedness, stability, and convergence of the numerical schemes, and discuss the related computational challenges. In the case of elliptic interface problems, we also present a posteriori error estimates.
@@ Introduction
The mathematical description of many physical and engineering problems requires partial differential equations involving different operators or discontinuous coefficients so that the computational domain is partitioned into several (possibly time dependent) regions. The coupling between different models is enforced at the interface via suitable transmission conditions. Among the applications, we mention, for instance, thermo-mechanical problems [1, 2], cardiac simulations [3, 4], computational geo-science [5, 6], and many others.
@@
When designing effective numerical methods for the mentioned problems, an accurate representation of the underlying geometry is needed. Fitted approaches are based on a mesh which conforms to the interface and provide accurate solutions [7, 8, 9, 10, 11]. However, when applied to evolution problems, computational issues may arise. For instance, the Arbitrary Lagrangian–Eulerian approach [12, 13, 14], appealing for fluid-structure interaction problems, needs an update of the mesh at each time step. In case of large displacements or deformations, the update may break shape regularity, thus requiring the generation of a new mesh, a costly operation in general. Moreover, when fluid and solid have similar or equal densities, the added mass effect may affect the performance of the method [15]. Such phenomenon may lead to the unconditional instability of the scheme regardless of the physical parameters. The added mass effect can be reduced by an appropriate treatment of transmission conditions, as described in [16, 17].
@@
Due to the critical aspects of fitted approaches, several unfitted methods have been introduced during the past decades. In this case, meshes are generated without taking care of the interface, which is allowed to cross elements. We mention, for instance, the Immersed Boundary Method [3, 18], where the interface position is tracked by Dirac delta functions, and the level set method [19, 20], where the interface corresponds to the zero level set of a certain function. Other techniques such as Nitsche-XFEM [21], Cut-FEM [22, 23, 24], Finite Cell methods [25, 26, 27], and Ghost-FEM [28] are based on the use of penalty terms. The fictitious domain approach has been introduced to model the particulate flow as described in [29, 30], while the application to more general cases can be found in [31, 32]. Another paradigm for tackling moving boundaries is given by diffuse interface approaches [33, 34, 35], where quantities localized on the interface are distributed throughout a thick interfacial region.
@@
In [36, 37], we introduced a new fictitious domain formulation with distributed Lagrange multiplier (FD-DLM) for the discretization of elliptic interface and fluid-structure interaction problems. Indeed, these problems share the common feature that the domain is divided into two regions separated by an interface. In the case of elliptic interface problems this is due to the presence of discontinuous coefficients, while fluid-structure interactions involve different types of equations, corresponding to constitutive models for fluid and solid. This survey paper aims to collect, in a unified framework, the main properties of this approach, including the well-posedness and stability of the solution, convergence of the discrete scheme and computational aspects.
@@
The main idea of our approach is to extend one domain into the other. The two domains are then discretized by fixed independent meshes. In particular, in case of fluid-structure interactions, the fluid dynamics is described in Eulerian framework on the extended domain, while the deformation of the solid body is represented in Lagrangian setting by considering a reference domain which is mapped, at each time instant, into the actual configuration. In order to impose that the solution in the fictitious region coincides with the solution in the immersed domain, we consider a Lagrange multiplier and a coupling term is added to the model. We emphasize that all computations are done on fixed domains, so that meshes are generated only once.
@@
The paper is organized into two parts. In Section 1, we focus on elliptic interface problems and discuss in some details the main features of our approach, which will be adapted later on for fluid-structure interactions. After presenting the continuous and discrete problems, we describe how to deal with the coupling term [38, 39]. Indeed, the construction of its finite element counterpart requires integration over non-matching grids, which is a challenging aspect of several immersed methods [40, 41, 42, 43, 44]. Our formulation turns out to be quite robust from this point of view since, in some situations, optimal results can be achieved even if the coupling term is computed in approximate way. Moreover, the presence of small cut cells does not affect the stability and conditioning. Next, we discuss the design of effective block preconditioners, whose action is performed thorough direct inversion [45] or by employing multigrid algorithms [10]. Another issue is related to the fact that the solution of the interface problem can present singularities along the interface. A way to achieve optimal convergence properties is to employ adaptive mesh refinement based on suitable error indicators. This crucial aspect has been previously analyzed in e.g. [46, 47, 48, 49] for other numerical schemes. Thus, we report the a posteriori error analysis presented in [50] for our FD-DLM formulation.
@@
The second section of the present paper extends the FD-DLM formulation and the related techniques to fluid-structure interaction problems, where an incompressible structure made of viscous hyper-elastic material is immersed in an incompressible Newtonian fluid (see [37] and references therein). In this case, the regions occupied by the fluid and the solid evolve in time and their configuration is itself an unknown of the problem. The actual position of the solid is obtained by mapping the reference domain through the deformation unknown: this gives further difficulties for the computation of the so-called coupling terms. Viscous hyper-elastic materials are useful when dealing with several applications: we mention, for instance, blood flow in heart [3, 51] and vessels [52, 53, 4], and flapping flexible filaments in soap films [54, 55]. This choice of solid constitutive law is relevant for our formulation since it allows to separate the viscous and elastic parts of the Cauchy stress tensor so that we can treat each of them in the appropriate Eulerian and Lagrangian frameworks, respectively.
@@
Both sections 1 and 2 are completed by some numerical results that confirm the theoretical findings and show the versatility of the proposed method. Moreover, several comments throughout this work are devoted to compare the main properties of our formulation with those of other unfitted techniques. For a complete and systematic comparison, we refer the reader to [56].
@@
Sobolev spaces are denoted by the standard symbol $W^{s,p}(D)$, where $s\in\mathbb{R}$ refers to differentiability, while $p\in[1,\infty]$ is the integrability exponent. If $p=2$, then the notation $H^{s}(D)=W^{s,2}(D)$ is employed, with associated norm $\|\cdot\|_{s,D}$ and semi-norm $|\cdot|_{s,D}$.
@@ 1. Fictitious domain formulation for elliptic interface problems
Let $\Omega\subset\mathbb{R}^{d}$, with $d=1,2,3$, be an open bounded domain with Lipschitz boundary $\partial\Omega$. The domain $\Omega$ is made of two disjoint regions, $\Omega_{1}$ and $\Omega_{2}$, separated by the interface $\Gamma=\overline{\Omega}_{1}\cap\overline{\Omega}_{2}$}. We assume $\Gamma$ to be Lipschitz continuous. Two examples of domain configuration are sketched in Figure 1.
@@
$-\operatorname{div}\left(\nu_{1}\operatorname{\nabla}u_{1}\right)$ $=f_{1}$ $\text{in }\Omega_{1}$
$-\operatorname{div}\left(\nu_{2}\operatorname{\nabla}u_{2}\right)$ $=f_{2}$ $\text{in }\Omega_{2}$
$u_{1}-u_{2}$ $=0$ $\text{on }\Gamma$
$\nu_{1}\,\nabla u_{1}\cdot\mathbf{n}_{1}+\nu_{2}\,\nabla u_{2}\cdot\mathbf{n}_{2}$ $=0$ $\text{on }\Gamma$
$u_{1}$ $=0$ $\text{on }\partial\Omega_{1}\setminus\Gamma$
$u_{2}$ $=0$ $\text{on }\partial\Omega_{2}\setminus\Gamma.$
@@
[…]
the variational formulation of Problem 1.1 can be written by applying standard arguments as follows:
@@
Equation (1) is equivalent to Problem 1.1 with $\tilde{u}_{|\Omega_{i}}=u_{i}$ with $i=1,2$.
Remark 1.2.
In general, due to the jump of the normal derivative of $\tilde{u}$ on $\Gamma$, the solution of (1) does not belong to $H^{2}(\Omega)$. Indeed, it is well-known that the regularity of the solution to elliptic interface problems with discontinuous coefficients on a domain with Lipschitz boundary belongs to $H^{r}(\Omega)$ with $1<r<3/2$. However, it is possible to show that the restriction $\tilde{u}_{|\Omega_{i}}$, $i=1,2$, is more regular depending on the smoothness of $\Gamma$. If $\Gamma$ is only Lipschitz continuous and presents re-entrant corners, then $u_{i}\in H^{s}(\Omega_{i})$ with $3/2<s\leq 2$ (see [57]).
Therefore, solving (1) by the finite element method on a mesh which does not take into account the presence of $\Gamma$, gives a non-optimal order of convergence. An optimal method can be recovered by constructing a mesh fitted with the interface $\Gamma$. Clearly, this requirement is strong, especially if the problem under consideration derives from the time discretization of a more complex one, such as a fluid-structure interaction problem. In such case, the interface $\Gamma$ may evolve in time and, thus the mesh should be updated according to the position of $\Gamma$ at each time step. Our approach is to use the fictitious domain formulation with distributed Lagrange multiplier.
@@ 1.1. Fictitious domain formulation
We denote by $u\in H^{1}_{0}(\Omega)$ the extension of $u_{1}$ to $\Omega$ satisfying $u_{|\Omega_{1}}=u_{1}$. The extended $u$ is required to match $u_{2}\in H^{1}(\Omega_{2})$ in $\Omega_{2}$. We enforce this superimposition between $u_{|\Omega_{2}}$ and $u_{2}$ at variational level by introducing a suitable functional space $\Lambda$ and a bilinear form $c:\Lambda\times H^{1}(\Omega_{2})\rightarrow\mathbb{R}$ satisfying
@@
$\int_{\Omega_{2}}(\varphi\cdot v_{2}+\operatorname{\nabla}\varphi\cdot\operatorname{\nabla}v_{2})\,dx_{2}=\langle\mu,v_{2}\rangle\qquad\forall v_{2}\in H^{1}(\Omega_{2})$
with $\|\varphi\|_{1,\Omega_{2}}=\|\mu\|_{\Lambda}$.
We introduce the Lagrange multiplier $\lambda\in\Lambda$ associated with the condition $u_{|\Omega_{2}}=u_{2}$, then the weak formulation of Problem 1.1 in a fictitious domain setting is:
Problem 1.3.
@@
In [36, Thm. 2], it has been proved that Problem 1.3 is equivalent to the standard variational formulation in (1). The following proposition states the well-posedness of Problem 1.3, see [36, Prop. 1].
Proposition 1.4.
Let $\nu$ and $\nu_{2}$ be two positive constants. Given $f\in L^{2}(\Omega)$ and $f_{2}\in L^{2}(\Omega_{2})$, Problem 1.3 admits a unique solution $(u,u_{2},\lambda)\in H^{1}_{0}(\Omega)\times H^{1}(\Omega_{2})\times\Lambda$} satisfying
@@
Since the problem shows a saddle point structure, its well-posedness is established by proving the following sufficient conditions [59]:
@@
Remark 1.5.
The stability estimate in Proposition 1.4 depends on the choice of $\Lambda$ and $c(\cdot,\cdot)$ through the constant $C$.
Remark 1.6.
Notice that, due to the equivalence of Problem 1.3 and (1), we have that the extended solution $u$ belongs to $H^{r}(\Omega)$ with $1<r<3/2$. On the other hand, $u_{2}$ belongs to $H^{s}(\Omega_{2})$ with $3/2<s\leq 2$, see Remark 1.2.
@@ 1.2. Discretization
We discretize Problem 1.3 by mixed finite elements. We partition $\Omega$ and $\Omega_{2}$ by two independent shape regular meshes $\mathcal{T}_{h}$ and $\mathcal{T}_{h}^{2}$, respectively. We denote by $h$ the size of $\mathcal{T}_{h}$, whereas $h_{2}$ denotes the mesh size of $\mathcal{T}_{h}^{2}$. The arguments we are going to present in this section are valid for both triangular and quadrilateral meshes, although we present in detail the quadrilateral case.
@@
On the other hand, if $\Lambda=H^{1}(\Omega_{2})$, the inner product in $H^{1}(\Omega_{2})$ can be used also at discrete level if $\Lambda_{h}$ contains continuous functions.
The discrete formulation of Problem 1.3 reads as follows.
Problem 1.7.
@@
For a generic element $K$, we denote by $\mathcal{P}_{1}(K)$ the space of linear polynomials if $K$ is a simplex and by $\mathcal{Q}_{1}(K)$ the space of bilinear polynomials if $K$ is a quadrilateral or an hexahedron. Moreover, let $\mathcal{P}_{0}(K)$ be the space of constants and $\mathfrak{B}(K)$ the space of bubbles vanishing at the boundary of an element $K$.
We solve Problem 1.7 by considering two distinct sets of finite element spaces, which are distinguished by the Lagrange multiplier space being continuous or discontinuous depending on the choice of the coupling term. More precisely, we consider continuous linear elements for all variables [36], that is
(6a) $V_{h}=\left\{v_{h}\in V:v_{h|K}\in\mathcal{Q}_{1}(K),\forall K\in\mathcal{T}_{h}\right\},$
(6b) $V_{2,h}=\left\{v_{2,h}\in V_{2}:v_{2,h|K_{2}}\in\mathcal{Q}_{1}(K_{2}),\forall K_{2}\in\mathcal{T}_{h}^{2}\right\},$
(6c) $\Lambda_{h}=V_{2,h}$
@@
(7a) $V_{h}=\left\{v_{h}\in V:v_{h|K}\in\mathcal{Q}_{1}(K),\forall K\in\mathcal{T}_{h}\right\}\subset V,$
(7b) $V_{2,h}=\left\{v_{2,h}\in V_{2}:v_{2,h|K_{2}}\in\mathcal{Q}_{1}(K_{2})\oplus\mathfrak{B}(K_{2}),\forall K_{2}\in\mathcal{T}_{h}^{2}\right\}\subset V_{2},$
(7c) $\Lambda_{h}=\left\{\lambda_{h}\in\Lambda:\lambda_{h|K_{2}}\in\mathcal{P}_{0}(K_{2}),\forall K_{2}\in\mathcal{T}_{h}^{2}\right\}\subset\Lambda.$
@@
Both choices are stable since they satisfy the discrete counterpart of (3) and (4). In particular, stability of (6) has been proved for both simplicial and quadrilateral/hexahedral meshes in [36, 58]. The ellipticity in the kernel condition is met under the assumption that $\nu_{2}>\nu$. This requirement can be relaxed for simplicial meshes if $h_{2}/h^{d}$ is small enough, see [58]. Stability of (7) has been proved in [60]. The ellipticity in the kernel condition is satisfied assuming that $\nu_{2}>\nu$. Numerical findings suggest that this requirement might be relaxed [60].
The error estimates on the discrete solution are provided in [36, 58, 60]. We point out that such bounds depend on the choice of $\Lambda_{h}$. The following estimate holds.
Proposition 1.8.
Let $(u,u_{2},\lambda)$ and $(u_{h},u_{2h},\lambda_{h})$ be the solutions of Problems 1.3 and 1.7, respectively. If the discrete spaces are given by (6), then the following error estimate holds true
@@
where $r$, $s$ are given in Remark 1.2, $1/2<t<1$ and $C_{1},\,C_{2}$ are positive constants independent of $h$ and $h_{2}$.